I have $14.80 in dimes and quarters the number of quarters is 2 less than 3tims the number of dimes.
How many quarters total?

Answer by Alex2dadeath(5)

Dimes will be x. A dime is worth 10 cents. So we will start off with .10x (the value of the dime and the x since we are looking for the number of dimes. Now it says there are 2 less than 3 times the number of dimes. Now as you see, that will be marked 3x (3 times the number) and 2 less which will be minus 2 (-2) so you will have 3x-2. A quarter is worth 25 cents so now we have to set up an equation also counting the values of the coins into it so you will now have: .10x + .25(3x-2) = 14.80. Start off by plugging in 25 inside parenthesis which you will then get: .10x + .75x - .50 = $14.80.
Now add .50 to both sides of the equation and also combine like terms (the x's add together) and now you will have: .85x = 15.30. Now this seems off since we are working with cents so you have .85 originally and 15.30. Let's move the decimal to the right on each side twice to the right to leave you with whole numbers to make it easier. So you now have: 85x = 1530. Divide both sides by 85 to leave x alone. x=18.
Now we put x into the top. 10x is 10(18) equals 18 dimes. And as for quarters, you got 3x minus 2 so now replace 18 for the x: 3(18)-2. 18x3= 54-2= 52 quarters.
Now double check your answer to make sure it adds up to the value. .10(dime amount) times the number of dimes (18) equals 1.80. subtract 1.80 from 14.18. You got $13 in quarters. Sum up the amount of quarters
$13 dollars in quarters is 52.

-ANSWER: 52 QUARTERS AND 18 DIMES-